# Language Models Encode Collaborative Signals in Recommendation

### v1

@@ 2 Uncovering Collaborative Signals in LMs via Linear Mapping
Table 1: The recommendation performance of linear mapping comparing with classical CF baselines.
Books Movies & TV Video Games
@@
CF MF (Rendle et al., 2012) $0.0437$ $0.0391$ $0.2476$ $0.0568$ $0.0519$ $0.3377$ $0.0323$ $0.0195$ $0.0864$
MultVAE (Liang et al., 2018) $0.0722$ $0.0597$ $\underline{0.3418}$ $0.0853$ $0.0776$ $0.4434$ $0.0908$ $0.0531$ $0.2211$
LightGCN (He et al., 2021) $0.0723$ 0.0608 0.3489 $0.0849$ $0.0747$ $0.4397$ $0.1007$ $0.0590$ $0.2281$
Linear Mapping BERT 0.0226 0.0194 0.1240 0.0415 0.0399 0.2362 0.0524 0.0309 0.1245
RoBERTa 0.0247 0.0209 0.1262 0.0406 0.0387 0.2277 0.0578 0.0338 0.1339
Llama2-7B 0.0662 0.0559 0.3176 0.1027 0.0955 0.4952 0.1249 0.0729 0.2746
Mistral-7B 0.0650 0.0544 0.3124 0.1039 0.0963 0.4994 0.1270 0.0687 0.2428
text-embedding-ada-v2 0.0515 0.0436 0.2570 0.0926 0.0874 0.4563 0.1176 0.0683 0.2579
text-embeddings-3-large 0.0735 0.0608 0.3355 0.1109 0.1023 0.5200 0.1367 0.0793 0.2928
SFR-Embedding-Mistral 0.0738 0.0610 0.3371 0.1152 0.1065 0.5327 0.1370 0.0787 0.2927
Table 1 reports the recommendation performance yielded by the linear mapping on three Amazon datasets [1], comparing with classic CF baselines. We observe that the performance of most advanced text embedding models (e.g., text-embeddings-3-large [73] and SFR-Embedding-Mistral [74]) exceeds LightGCN on all datasets. We also empirically show that these improvements may not merely come from the better feature encoding ability (refer to Appendix B.3). Figure 1(c) further suggests that obscure user preference similarities may be implicitly encoded in the language representation space, with a linear mapping matrix to activate. These findings indicate the homomorphism [54] between the language representation spaces of advanced LMs and effective item representation spaces for recommendation. In addition, with the help of linear mapping, for the first time, we are able to clearly quantify the gap in the representation spaces between different LMs for recommendation. With the advances in LMs, we observe that the performance of language representations exhibits a rising trend, especially showing a sharp rise from BERT-style models to LLMs. Representations from early BERT-style models (e.g., BERT [4] and RoBERTa [5]) only show weaker or equal capabilities compared with MF, failing behind LLMs largely. This further suggests that, until very recently, the language representation spaces are able to encode user preference similarities efficiently.
3 AlphaRec
This finding of space homomorphism [54] sheds light on building advanced CF models purely based on language representations without introducing ID-based embeddings. Inspired by this, we develop a simple yet effective CF model called AlphaRec, by incorporating only three simple components (i.e., nonlinear projection [62], graph convolution [56] and contrastive learning (CL) objectives [57]). It is important to highlight that our approach is centered on exploring the potential of language representations for recommendation by integrating essential components from leading CF models, rather than deliberately inventing new CF mechanisms. Surprisingly, these basic and simple components lead to an excellent language-representation-based recommender with multiple capabilities. We present the model structure of AlphaRec in Section 3.1, and compare AlphaRec with two popular recommendation paradigms in Section 3.2.
3.1 Method
@@
We present how AlphaRec is designed and trained. Generally, the representation learning architecture $\phi_{\theta}(\cdot,\cdot)$ of AlphaRec is simple, which only contains a two-layer MLP and the basic graph convolution operation, with language representations as the input features $\mathbf{x}_{u},\mathbf{x}_{i}$. The cosine similarity is used as the similarity function $s(\cdot,\cdot)$, and the contrastive loss InfoNCE [57, 58] is adopted for optimization. For simplicity, we consistently adopt text-embeddings-3-large [73] as the language representation model, for its excellent language understanding and representation capabilities.
@@
Nonlinear projection. In AlphaRec, we substitute the linear mapping matrix delineated in Section 2 with a nonlinear MLP. This conversion from linear to nonlinear is non-trivial, for the paradigm shift from ID-based embeddings to language representations, since nonlinear transformation helps in excavating more comprehensive user preference similarities from the language representation space with rich semantics (see discussions about this in Appendix C.2.3) [62]. Specifically, we project the language representation $\mathbf{x}_{i}$ of the item title to an item space for recommendation with the two-layer MLP, and obtain user representations as the average of historical items:
$\mathbf{e}_{i}^{(0)}=\bm{W}_{2}\operatorname{LeakyReLU}\left(\bm{W}_{1}\mathbf{x}_{i}+\bm{b}_{1}\right)+\bm{b}_{2},\quad\mathbf{e}_{u}^{(0)}=\frac{1}{|\mathcal{N}_{u}|}\sum_{i\in\mathcal{N}_{u}}\mathbf{e}^{(0)}_{i}.$ (1)
@@
Graph convolution. Graph neural networks (GNNs) have shown superior effectiveness for recommendation [52, 56], owing to the natural user-item graph structure in recommender systems [75]. In AlphaRec, we employ a minimal graph convolution operation [56] to capture more complicated collaborative signals from high-order connectivity [56, 76, 77, 75] as follows:
@@
The information of connected neighbors is aggregated with a symmetric normalization term $\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}\sqrt{\left|\mathcal{N}_{i}\right|}}$. Here $\mathcal{N}_{u}$ ($\mathcal{N}_{i}$) denotes the historical item (user) set that user $u$ (item $i$) has interacted with. The features $\mathbf{e}^{(0)}_{u}$ and $\mathbf{e}^{(0)}_{i}$ projected from the MLP are used as the input of the first layer. After propagating for $K$ layers, the final representation of a user (item) is obtained as the average of features from each layer:
@@
Contrastive learning objective. The introduction of contrasting learning is another key element for the success of leading CF models. Recent research suggests that the contrast learning objective, rather than data augmentation, plays a more significant role in improving recommendation performance [70, 78, 69]. Therefore, we simply use the contrast learning object InfoNCE [57] as the loss function without any additional data augmentation on the graph [79, 58]. With cosine similarity as the similarity function $s(\mathbf{e}_{u},\mathbf{e}_{i})=\frac{{\mathbf{e}_{u}}^{\top}\mathbf{e}_{i}}{\left\lVert\mathbf{e}_{u}\right\rVert\cdot\left\lVert\mathbf{e}_{i}\right\rVert}$, the InfoNCE loss [57, 79, 80] is written as:
@@
Here, $\tau$ is a hyperparameter called temperature [81], $\mathcal{O}^{+}=\{(u,i)|y_{ui}=1\}$ denoting the observed interactions between users $\mathcal{U}$ and items $\mathcal{I}$. And $\mathcal{S}_{u}$ is a randomly sampled subset of negative items that user $u$ does not adopt.
3.2 Discussion of Recommendation Paradigms
We compare the language-representation-based AlphaRec with two popular recommendation paradigms in Table 2 (see more discussion about related works in Appendix A).
ID-based recommendation (ID-Rec) [52, 55]. In the traditional ID-based recommendation paradigm, users and items are represented by ID-based learnable embeddings derived from a large number of user interactions. While ID-Rec exhibits excellent recommendation capabilities with low training and inference costs [63, 79], it also has two significant drawbacks. Firstly, these ID-based embeddings learned in specific domains are difficult to transfer to new domains without overlapping users and items [37], thereby hindering zero-shot recommendation capabilities. Additionally, there is a substantial gap between ID-Rec and natural languages [34], which makes ID-based recommenders hard to incorporate language-based user intentions and further refine recommendations accordingly.
LM-based recommendation (LM-Rec) [15, 16, 24]. Benefitting from the extensive world knowledge and powerful reasoning capabilities of LMs [7, 82], the LM-based recommendation paradigm has gained widespread attention [11, 13]. LM-Rec tends to convert user interaction history into text prompts as input for LMs, utilizing pre-trained or fine-tuned LMs in a text generation pattern to recommend items. LM-Rec demonstrates zero-shot and few-shot abilities and can easily understand language-based user intentions. However, LM-Rec faces significant challenges. Firstly, the LM-based model architecture leads to huge training and inference costs, with real-world deployment difficulties. Additionally, limited by the text generation paradigm, LM-based models tend to perform candidate selection [17] or generate a single next item [24]. It remains difficult for LM-Rec to comprehensively rank the entire item corpus or recommend multiple items that align with user interests.
Language-representation-based recommendation. We argue that AlphaRec follows a new CF paradigm, which we term the language-representation-based paradigm. This paradigm replaces the ID-based embeddings in ID-Rec with representations from pre-trained LMs, employing feature encoders to map language representations directly into the recommendation space. Few early studies lie in this paradigm, including using BERT-style LMs to learn universal sequence representations [37, 44], or adopting the same model architecture as ID-Rec with simple input features replacement [34, 35]. These early explorations, which are mostly based on BERT-style LMs, are usually only applicable in certain specific scenarios, such as the transductive setting with the help of ID-based embeddings [37]. This phenomenon is consistent with our previous findings in Section 2, indicating that BERT-style LMs may fail to effectively encode user preference similarities. We point out that AlphaRec is the first recommender in the language-representation-based paradigm to surpass the traditional ID-based paradigm on multiple tasks, faithfully demonstrating the effectiveness and potential of this paradigm.
Table 2: Comparison of recommendation paradigms
Recommendation Paradigms Training Cost Zero-shot Ability Intention-aware Ability
ID-based Low
LM-based High
Language-representation-based Low
4 Experiments
In this section, we aim to explore the effectiveness of AlphaRec. Specifically, we are trying to answer the following research questions:
  • RQ1: How does AlphaRec perform compared with leading ID-based CF methods?
  • RQ2: Can AlphaRec learn general item representations, and achieve good zero-shot recommendation performance on entirely new datasets?
  • RQ3: Can AlphaRec capture user intention described in natural language and adjust the recommendation results accordingly?
4.1 General Recommendation Performance (RQ1)
Motivation. We aim to explore whether the language-representation-based recommendation paradigm can outperform the ID-Rec paradigm. An excellent performance of AlphaRec would shed light on the research line of building language-representation-based recommenders in the future.
@@ 4.1 General Recommendation Performance (RQ1)
Baselines. We consider ID-based baselines in this section. LM-based methods are not incorporated for two practical reasons: the huge inference cost on datasets with millions of interactions and the task limitation of candidate selection or next item prediction. In addition to classic baselines (i.e., MF, MultVAE, and LightGCN) introduced in section 2, we consider two categories of leading ID-based CF baselines: CL-based CF methods: SGL [83], BC Loss [79], XSimGCL [70] and LM-enhanced CF methods: KAR [48], RLMRec [47]. See more details about baselines in Appendix C.2.1.
@@
Results. Table 3 presents the performance of AlphaRec compared with leading CF baselines. The best-performing methods are bold, while the second-best methods are underlined. Figure 2(a) and Figure 2(b) report the training efficiency and ablation results. We observe that:
@@
Books Movies & TV Video Games
@@
MF (Rendle et al., 2012) $0.0437$ $0.0391$ $0.2476$ $0.0568$ $0.0519$ $0.3377$ $0.0323$ $0.0195$ $0.0864$
MultVAE (Liang et al., 2018) $0.0722$ $0.0597$ $0.3418$ $0.0853$ $0.0776$ $0.4434$ $0.0908$ $0.0531$ $0.2211$
LightGCN (He et al., 2021) $0.0723$ $0.0608$ $0.3489$ $0.0849$ $0.0747$ $0.4397$ $0.1007$ $0.0590$ $0.2281$
SGL (Wu et al., 2021) 0.0789 0.0657 0.3734 0.0916 0.0838 0.4680 0.1089 0.0634 0.2449
BC Loss (Zhang et al., 2022) 0.0915 0.0779 0.4045 0.1039 0.0943 0.5037 0.1145 0.0668 0.2561
XSimGCL (Yu et al., 2024) 0.0879 0.0745 0.3918 0.1057 0.0984 0.5128 0.1138 0.0662 0.2550
KAR (Xi et al., 2023) 0.0852 0.0734 0.3834 0.1084 0.1001 0.5134 0.1181 0.0693 0.2571
RLMRec (Ren et al., 2024a) 0.0928 0.0774 0.4092 0.1119 0.1013 0.5301 0.1384 0.0809 0.2997
AlphaRec 0.0991* 0.0828* 0.4185* 0.1221* 0.1144* 0.5587* 0.1519* 0.0894* 0.3207*
@@
(a) Training efficiency comparison
(b) Ablation study on Books
@@
Figure 2: (2(a)) The bar charts show the number of epochs needed for each model to converge. AlphaRec tends to exhibit an extremely fast convergence speed. (2(b)) The effect of each component in AlphaRec on Books dataset.
@@
  • AlphaRec consistently outperforms leading CF baselines by a large margin across all metrics on all datasets. AlphaRec shows an improvement ranging from 6.79% to 9.75% on Recall@20 compared to the best baseline RLMRec [47]. We further conduct the ablation study to explore the reason for its success (see more ablation results in Appendix C.2.2). As shown in Figure 2(b), each component in AlphaRec contributes positively. Specifically, the performance degradation caused by replacing the MLP with a linear weight matrix (w/o MLP) indicates that nonlinear transformations can extract the implicit user preference similarities encoded in the language representation space more effectively. Moreover, the performance drop from replacing InfoNCE loss [58] with BPR loss [71] (w/o CL) and removing the graph convolution (w/o GCN) suggests that explicitly modeling the collaborative relationships through the loss function and model architecture can further enhance recommendation performance. These findings suggest that, by carefully designing the model to extract user preference similarities, the potential of language-representation-based paradigm can be realized, surpassing the leading ID-Rec paradigm.
@@
  • The incorporation of semantic language representations into traditional ID-based CF methods can lead to significant performance improvements. We note that two LM-enhanced CF methods, KAR and RLMRec, both show improvements over CL-based CF methods. Nevertheless, the combination of ID-based embeddings and language representations in these methods does not yield higher results than purely language-representation-based AlphaRec. We attribute this phenomenon to the fact that the performance contribution of these methods mainly comes from the language representations, which is consistent with the previous findings [34, 44].
  • AlphaRec exhibits fast convergence speed. We find that the convergence speed of AlphaRec is comparable with, or even surpasses, CL-based methods with data augmentation (e.g., SGL [83] and XSimGCL [70]). Meanwhile, methods based solely on graph convolution (LightGCN [56]) or CL objective (BC Loss [79]) show relatively slow convergence speed, indicating that introducing these modules may not lead to convergence speed improvement. We attribute the fast convergence speed of AlphaRec to the homomorphism between the language space and a good recommendation space. As a result, only minor adjustments to the vanilla language representations are needed for effective recommendation.
4.2 Zero-shot Recommendation Performance on Entirely New Datasets (RQ2)
Motivation. We aim to explore whether AlphaRec has learned general item representations [37], which enables it to perform well on entirely new datasets without any user and item overlap.

### v2

@@ 3.2 Empirical Findings
Table 1: The comparison of the recommendation performance of linear mapping with the classical ID-based CF baselines.
Movies & TV Video Games Books
@@
CF MF (Rendle et al., 2012) $0.0568$ $0.0519$ $0.3377$ $0.0323$ $0.0195$ $0.0864$ $0.0437$ $0.0391$ $0.2476$
MultVAE (Liang et al., 2018) $0.0853$ $0.0776$ $0.4434$ $0.0908$ $0.0531$ $0.2211$ $0.0722$ $0.0597$ $\underline{0.3418}$
LightGCN (He et al., 2021) $0.0849$ $0.0747$ $0.4397$ $0.1007$ $0.0590$ $0.2281$ $0.0723$ 0.0608 0.3489
Linear Mapping BERT 0.0415 0.0399 0.2362 0.0524 0.0309 0.1245 0.0226 0.0194 0.1240
RoBERTa 0.0406 0.0387 0.2277 0.0578 0.0338 0.1339 0.0247 0.0209 0.1262
Llama2-7B 0.1027 0.0955 0.4952 0.1249 0.0729 0.2746 0.0662 0.0559 0.3176
Mistral-7B 0.1039 0.0963 0.4994 0.1270 0.0687 0.2428 0.0650 0.0544 0.3124
text-embedding-ada-v2 0.0926 0.0874 0.4563 0.1176 0.0683 0.2579 0.0515 0.0436 0.2570
text-embeddings-3-large 0.1109 0.1023 0.5200 0.1367 0.0793 0.2928 0.0735 0.0608 0.3355
SFR-Embedding-Mistral 0.1152 0.1065 0.5327 0.1370 0.0787 0.2927 0.0738 0.0610 0.3371
(a) Movie & TV
(b) Games
(c) Books
Figure 2: The recommendation performance of linear mapping with different language model sizes.
Existence (RQ1). To explore the existence of collaborative signals in language representations, we test the recommendation performance of the linear mapping method. Table 1 reports the performance yielded by post-mapping representations on three Amazon datasets (Ni et al., 2019), comparing with classic ID-based CF baselines: matrix factorization (MF) (Koren et al., 2009), MultVAE (Liang et al., 2018), and LightGCN (He et al., 2021) (see baseline details in Appendix C.2). Figures (2(a)) - (2(c)) depict the linear mapping performance under different LM sizes. Figure 1(c) demonstrates the visualization of representations before and after linear mapping. We observe that:
  • Post-mapping representations of advanced LMs achieve superior recommendation performance in most cases, suggesting the possible homomorphism between language spaces and behavior spaces. Specifically, advanced LMs (e.g., Llama2-7B (Touvron et al., 2023b) and text-embeddings-3-large (Neelakantan et al., 2022)) consistently perform better than leading CF models (e.g., LightGCN) on most metrics. We also see that the performance improves with more recent and advanced LMs. In contrast, earlier BERT-style models (e.g., BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019)) perform similarly to or worse than MF, indicating that LMs have only recently developed the ability to encode user preference similarities effectively.
  • Language representations encode user preference similarities beyond semantic textual similarities (STS). Consider Figure 1(c) as an example again, homosexual movies, which differ significantly in textual meaning, cluster together after linear mapping. This suggests that user preferences, which are not immediately apparent from text alone, are implicitly encoded in the language space and can be uncovered through linear mapping.
Table 2: The robustness of language representations for recommendation.
Movies & TV Video Games Books
Recall NDCG HR Recall NDCG HR Recall NDCG HR
Title + Random Noise 0.0952 0.0887 0.4731 0.1213 0.0706 0.2722 0.0632 0.0525 0.3099
Title Only 0.1027 0.0955 0.4952 0.1249 0.0729 0.2746 0.0662 0.0559 0.3176
Scaling and Robustness (RQ2). We also investigate whether the encoding of user preference similarities in LMs scales with model size and whether this encoding is robust in the presence of noise in the input prompts. To this end, we test the linear mapping performance of LMs of various sizes and prompts with noise:
  • The encoding of user preference similarities becomes more refined as model size increases, leading to better linear mapping performance. Specifically, we test the linear mapping performance across different language model sizes (7B, 13B, and 70B) of the Llama2 family (Touvron et al., 2023b). Llama3 (Dubey et al., 2024) is not selected due to the lack of a 13B model. As shown in Figure 2, linear mapping performance improves consistently as the model size increases from 7B to 70B, indicating that larger models capture more nuanced user behavioral patterns.
  • Language representations are relatively robust to prompt disturbances. Following previous works (Gurnee & Tegmark, 2023), we compare two prompting strategies: using item titles alone (e.g., Castlevania), and adding 5-10 random letters to the titles (e.g., Castlevania sdfhdsk). Table 2 shows that adding random noise to the item titles had minimal impact on the linear mapping performance. The prompt noise has more impact on Movies & TV since item titles in this dataset are relatively shorter than others (see Appendix C.1). This finding suggests the relative robustness of the recommendation knowledge encoded in the language representation space.
4 Leveraging Language Representations for Better Recommendation
This finding of possible space homomorphism (Dieudonne, 1969) and encoded collaborative signals arouse interest in the following questions. RQ3: How powerful are such language representations for building advanced CF models that can outperform prevailing ID-based CF methods? To address these questions, in Section 4.1, we aim to develop a simple yet effective CF model termed AlphaRec, which is solely based on language representations and merely incorporates three crucial components in modern CF models. After that, we evaluate its performance in Section 4.2 to demonstrate the capability of advanced language representations for recommendation.
4.1 AlphaRec
@@
We briefly present how this simple model AlphaRec is designed and trained. It is important to highlight that we center on exploring the power of language representations for recommendation, rather than deliberately inventing new CF mechanisms. Generally, the representation generation architecture $\phi_{\theta}(\cdot,\cdot)$ is simple, which only contains a two-layer MLP and the basic graph convolution operation. The cosine similarity is used as the similarity function $s(\cdot,\cdot)$, and the contrastive loss InfoNCE (van den Oord et al., 2018; Wu et al., 2022) is adopted for optimization. For simplicity, we adopt text-embeddings-3-large (Neelakantan et al., 2022) for language representation generation by default, for its excellent language understanding and representation capabilities.
@@
Nonlinear projection. We substitute the linear matrix delineated in Section 3 with a nonlinear MLP. Nonlinear transformation helps in excavating more comprehensive preference similarities from the language representation space (see discussions about this in Appendix C.4) (He et al., 2017). Taking the averaged language representaions of historical items as the user language representation (i.e., $\mathbf{z}_{u}=\frac{1}{|\mathcal{N}_{u}|}\sum_{i\in\mathcal{N}_{u}}\mathbf{z}_{i}$), the initial nonlinear transformation operation be formulated as:
$\mathbf{e}_{i}^{(0)}=\bm{W}_{2}\operatorname{LeakyReLU}\left(\bm{W}_{1}\mathbf{z}_{i}+\bm{b}_{1}\right)+\bm{b}_{2},\quad\mathbf{e}_{u}^{(0)}=\bm{W}_{2}\operatorname{LeakyReLU}\left(\bm{W}_{1}\mathbf{z}_{u}+\bm{b}_{1}\right)+\bm{b}_{2}.$ (1)
@@
Graph convolution. Graph neural networks (GNNs) show superior effectiveness for recommendation (Wang et al., 2019b), owing to the natural user-item graph structure in recommender systems (Wu et al., 2023b). We employ a minimal graph convolution operation (He et al., 2021) to capture more complicated collaborative patterns from high-order connectivity (Wu et al., 2019) as follows:
@@
The information of connected neighbors is aggregated with a symmetric normalization term $\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}\sqrt{\left|\mathcal{N}_{i}\right|}}$. Here $\mathcal{N}_{u}$ ($\mathcal{N}_{i}$) denotes the historical item (user) set that user $u$ (item $i$) has interacted with. The representations $\mathbf{e}^{(0)}_{u}$ and $\mathbf{e}^{(0)}_{i}$ projected from the MLP are used as the input of the first layer. After propagating for $K$ layers, the final behavior representation of a user $u$ (item $i$) is obtained as the average of representations from each layer:
@@
Contrastive learning objective. The introduction of contrastive learning (Radford et al., 2021) is another key element for the success of leading CF models. Recent research suggests that the contrast learning objective, rather than data augmentation, plays a more significant role in improving recommendation performance (Yu et al., 2024; 2022; Zhang et al., 2023a). Therefore, we simply use the contrast learning object InfoNCE (van den Oord et al., 2018) as the loss function without any additional data augmentation on the graph (Wu et al., 2022). With cosine similarity as the similarity function $s(\mathbf{e}_{u},\mathbf{e}_{i})=\frac{{\mathbf{e}_{u}}^{\top}\mathbf{e}_{i}}{\left\lVert\mathbf{e}_{u}\right\rVert\cdot\left\lVert\mathbf{e}_{i}\right\rVert}$, the InfoNCE loss (van den Oord et al., 2018) is written as:
@@
Here, $\tau$ is a hyperparameter called temperature (Wang & Liu, 2021), $\mathcal{O}^{+}=\{(u,i)|y_{ui}=1\}$ denoting the observed interactions between users $\mathcal{U}$ and items $\mathcal{I}$. And $\mathcal{S}_{u}$ is a randomly sampled subset of negative items that user $u$ does not adopt.
4.2 Empirical Findings
@@
Baselines. We compare AlphaRec with leading ID-based CF baselines, to assess the effectiveness of adopting advanced language representations. We do not consider baselines using LMs as recommenders for two practical reasons: the huge inference cost on datasets with millions of interactions and the task limitation of candidate selection (Liao et al., 2024) or next item prediction (Zheng et al., 2023). In addition to classic baselines introduced in section 3.2, we consider two categories of leading ID-based CF baselines, CL-based CF methods: SGL (Wu et al., 2021), BC Loss (Zhang et al., 2022), XSimGCL (Yu et al., 2024) and LM-enhanced CF methods: KAR (Xi et al., 2023), RLMRec (Ren et al., 2024b). See more details about baselines in Appendix C.2.
@@
Recommendation capabilities (RQ3). Table 3 presents performance comparison. The best-performing methods are bold, while the second-best methods are underlined. We observe that:
@@
Movies & TV Video Games Books
@@
MF (Rendle et al., 2012) $0.0568$ $0.0519$ $0.3377$ $0.0323$ $0.0195$ $0.0864$ $0.0437$ $0.0391$ $0.2476$
MultVAE (Liang et al., 2018) $0.0853$ $0.0776$ $0.4434$ $0.0908$ $0.0531$ $0.2211$ $0.0722$ $0.0597$ $0.3418$
LightGCN (He et al., 2021) $0.0849$ $0.0747$ $0.4397$ $0.1007$ $0.0590$ $0.2281$ $0.0723$ $0.0608$ $0.3489$
SGL (Wu et al., 2021) 0.0916 0.0838 0.4680 0.1089 0.0634 0.2449 0.0789 0.0657 0.3734
BC Loss (Zhang et al., 2022) 0.1039 0.0943 0.5037 0.1145 0.0668 0.2561 0.0915 0.0779 0.4045
XSimGCL (Yu et al., 2024) 0.1057 0.0984 0.5128 0.1138 0.0662 0.2550 0.0879 0.0745 0.3918
KAR (Xi et al., 2023) 0.1084 0.1001 0.5134 0.1181 0.0693 0.2571 0.0852 0.0734 0.3834
RLMRec (Ren et al., 2024b) 0.1119 0.1013 0.5301 0.1384 0.0809 0.2997 0.0928 0.0774 0.4092
AlphaRec 0.1221* 0.1144* 0.5587* 0.1519* 0.0894* 0.3207* 0.0991* 0.0828* 0.4185*
@@
(a) Ablation study on Books
(b) Training efficiency comparison
@@
Figure 3: (3(a)) The effect of each component on Books dataset. (3(b)) The number of epochs needed for each model to converge. AlphaRec exhibits a breakneck convergence speed.
@@
  • Advanced language representations shows strong potentials for recommendation, which can be unleashed by appropriate model design. AlphaRec consistently outperforms leading CF baselines by a large margin across all metrics on all datasets, with an improvement ranging from 6.79% to 9.75% on Recall@20 compared to the best baseline. Moreover, as shown in Figure 3(a), each component contributes positively (see more ablation results in Appendix C.3) in unleashing the power of language representations. Specifically, the performance degradation caused by replacing the MLP with a linear weight matrix (w/o MLP) indicates that nonlinear transformations can extract the implicit user preference similarities encoded in the language representation space more effectively. Besides, the performance also drops from replacing InfoNCE loss (Wu et al., 2022) with BPR loss (Rendle et al., 2012) (w/o CL) and removing the graph convolution (w/o GCN) suggests that explicitly modeling the collaborative relationships through the loss function and model architecture can further enhance recommendation performance. These findings suggest that the power of advanced language representations can be unleashed by carefully designing the model, showcasing the potential to surpass prevailing ID-based recommenders.
@@
  • The incorporation of advanced language representations can benefit traditional ID-based CF methods. We note that two LM-enhanced CF methods, KAR and RLMRec, both show improvements over the most advanced CF methods. Nevertheless, the combination of ID-based embeddings and language representations in these methods does not yield higher results than purely language-representation-based AlphaRec. We attribute this phenomenon to their naive design for the combination ID-based embeddings and language representations, which is also highlighted by previous works (Yuan et al., 2023; Zhang et al., 2024c).
5 Exploring Potentials of Language Representations for Recommendation
In this section, we focus on this question: What new opportunities beyond good performance can advanced language representations bring to recommender systems? To answer this, we systematically analyze AlphaRec and discover the following potentials of adopting such language representations. Potential 1: Good initialization for item representations (Section 5.1). Potential 2: Zero-shot ability (Section 5.2). Potential 3: Intention-aware ability (Section 5.3).
5.1 Good Initialization for Item Representations (Potential 1)
Advanced language representations may provide a good initialization for item representations, with few adjustments for effective recommendation. As shown in Figure 3(b), beyond its good performance, AlphaRec also exhibits extremely fast convergence speed, which is comparable with or even surpasses the fastest ID-based CF methods (e.g., SGL (Wu et al., 2021) and XSimGCL (Yu et al., 2024)). Moreover, recent works also suggest that, when using advanced language representations to initialize ID-based item embeddings (Harte et al., 2023; Zhao et al., 2024), the performance of traditional ID-based recommenders improves significantly. We attribute the above findings to the homomorphism between the language space and a good behavior space. Therefore, when using advanced language representations for initialization, only minor adjustments are needed to generate effective behavior representations for recommendation.
5.2 Zero-shot Ability (Potential 2)
The prevailing ID-based recommenders suffer from domain transferring problems (i.e., behavior representations are highly bound with ID information (Zhu et al., 2021)). Advanced language representations may provide opportunities for learning transferable item representations (Hou et al., 2022), enabling recommenders to perform well on entirely new datasets without any ID overlap. To address this potential, we test the zero-shot recommendation ability of AlphaRec (Ding et al., 2021).